Find the mean, variance, and standard deviation of the scores of the students in an English test: 8, 9, 12, 18, 11, 16, 25, and 19


1
Expert's answer
2022-02-11T12:49:05-0500
"mean=\mu"

"=\dfrac{8+9+12+18+11+16+25+19}{8}=14.75"

"Var(X)=\sigma^2=\dfrac{1}{8}((8-14.75)^2+(9-14.75)^2"

"+(12-14.75)^2+(18-14.75)^2+(11-14.75)^2"

"+(11-14.75)^2+(16-14.75)^2+(25-14.75)^2"

"+(19-14.75)^2)=29.4375"

"\sigma=\sqrt{\sigma^2}=\sqrt{29.4375}=5.4256"